# StPR: Spatiotemporal Preservation and Routing for Exemplar-Free Video Class-Incremental Learning

### v1

@@ Appendix C Additional Experiments
C.1 Complexity Analysis: FLOPs and Parameters
We compute the theoretical floating point operations (FLOPs) for each module: the CLIP ViT-B/16 backbone, the inserted adapters, and the spatiotemporal encoder module.
Table 7: FLOPs and parameter counts for each module per 8-frame video.
Module Input Shape Parameters FLOPs (GFLOPs)
CLIP (ViT-B/16) (frozen) $(8,3,224,224)$ 86M 269.81
Adapter $(8,197,768)$ 1.17M (1.36%) 3.73 (1.38%)
Spatiotemporal Encoder $(9,512)$ 9.45M (10.99%) 0.0854 (0.03%)
Table 7 provides a detailed breakdown of computational complexity and parameter count for each component in our framework, evaluated on 8-frame video inputs. The backbone CLIP (ViT-B/16) dominates the overall cost with 86M parameters and 148 GFLOPs. The inserted adapter modules, despite being integrated into every transformer layer, introduce only 1.17M additional parameters (1.36%) and 1.38% more FLOPs, demonstrating their lightweight nature. Furthermore, our spatiotemporal encoder—used to capture dynamic information—adds merely 0.085 GFLOPs (0.03%, per-expert) and 9.45M parameters (10.99%), confirming its computational efficiency. These results validate that our method enhances temporal modeling with minimal overhead, making it well-suited for continual learning in resource-constrained settings.
@@ C.2 Comparison of MoE Routing Strategies (task by task)
Figure 4 presents the relative accuracy change (%) of different Mixture-of-Experts (MoE) routing strategies on three benchmarks: UCF101 with two task configurations (5$\times$10s and 10$\times$5s), and HMDB51 (5$\times$5s). The horizontal axis represents the incremental task ID, while the vertical axis shows the accuracy change relative to the Avg-MoE baseline.
@@ C.3 Effectiveness of Frame-Shared Semantics Distillation (task by task)
Appendix D Limitations
One limitation of our approach is that the text classifier is based on a vision-language model pre-trained on natural images and general text corpora. This may introduce a domain gap when applied to specialized video domains (e.g., medical or industrial settings), where the semantic alignment between text and video content may be less reliable. Adapting or fine-tuning the text encoder on domain-specific data is a promising direction for future work.
Appendix E Potential Impact
Our method for video class-incremental learning has potential benefits in areas such as robotics and assistive systems by improving adaptability and efficiency, we acknowledge that any AI-based video understanding system could also be misused for surveillance or privacy-infringing applications. Mitigating such risks requires careful deployment and adherence to ethical standards.

### v2

@@ Appendix C Additional Experiments
C.1 Analysis of hyper-parameter.
Dataset Metric $\times 10^{3} 1×1$$\times 10^{4}$ $2.5\times 10^{4}$ $5\times 10^{4}$ $1\times 10^{5}$
HMDB($5\times 5$s) Acc $\uparrow$ 59.51 63.04 62.10 62.29 63.05
BWF $\downarrow$ 16.90 11.04 11.14 12.32 12.77
UCF101($5\times 10$s) Acc $\uparrow$ 85.10 85.79 84.54 84.88 84.16
BWF $\downarrow$ 7.93 5.63 6.95 5.67 6.83
Table. 7 analyzes the sensitivity of the hyper-parameter $w$, which controls the strength of FSSD regularization. Across a wide range of values, our framework maintains stable performance, indicating robustness to hyper-parameter variations. Moderate $w$ values achieve the best trade-off between knowledge retention and adaptability, avoiding under-regularization or excessive constraint on model plasticity.
@@ C.2 Comparison of MoE Routing Strategies (task by task)
[…]
@@
Figure 5: Heatmap of task selection by temporal decomposition-based router on the UCF101.
As shown in 5, temporal decomposition-based router is effective at task boundary decisions.
As shown in Figure 7, it maintains highest accuracy over time and effectively mitigates forgetting, demonstrating the complementary strengths of spatial semantic preservation and temporal dynamic modeling
@@ C.3 Effectiveness of Frame-Shared Semantics Distillation (task by task)
C.4 Task-by-task Ablation Study Visualization
As shown in Figure 7, we progressively add our modules (FSSD and TD-MOE) on the UCF101 and HMDB51 datasets. Significant improvements are observed at each incremental stage, with more pronounced gains as the number of tasks increases, especially in the long-term scenario (10 tasks). This further validates the effectiveness of our proposed method and the superior performance of our model.